OpenStudy (anonymous):

twice a number added to 3 is divded bythe number plus 8= seven fourths what is the number. I need help with this

OpenStudy (skullpatrol):

Step 1 Read the problem carefully. Decide what unknown numbers are asked for and what facts are known.

OpenStudy (skullpatrol):

Let n = the number

OpenStudy (anonymous):

well it seems ike it would be 2n+3/n+8+7?4

OpenStudy (skullpatrol):

What facts are known?

OpenStudy (anonymous):

twice a Number(n) added to 3 is divided b the number(n) plus 8 = 7/4

OpenStudy (skullpatrol):

Very good :-D

OpenStudy (skullpatrol):

Now translate your words into symbols.

OpenStudy (anonymous):

2n3/n+8=7/4

OpenStudy (skullpatrol):